# Modified gravity interpretation of the evolving dark energy in light of DESI data

### v1

@@ Abstract
Taking this preference seriously, we study a general Horndeski scalar-tensor theory, which provides a physical mechanism to safely cross the phantom divide, $w=-1$. Utilizing the Effective Field Theory of dark energy and adopting the $w_{0}w_{a}{\rm CDM}$ background cosmological model, we derive constraints on the parameters $w_{0}=-0.856\pm 0.062$ and $w_{a}=-0.53_{-0.26}^{+0.28}$ at $68\%$ CL from Planck CMB, Planck and ACT CMB lensing, DESI BAO, and Pantheon+ datasets, showing good consistency with the standard $w_{0}w_{a}{\rm CDM}$ model. The modified gravity model shows a preference over $\Lambda$CDM at the $2.4\sigma$ level, while for $w_{0}w_{a}{\rm CDM}$ it is at $2.5\sigma$. We conclude that modified gravity offers a viable physical explanation for DESI’s preference for evolving dark energy.
@@ I Introduction
One particular aspect of the preference of the DESI+CMB+SN data for an evolving dark energy equation of state is that the preferred models cross the ‘phantom divide’, $w=-1$, which was also noticed by the DESI collaboration DESI:2024kob . To some extent this is driven by the CMB data, which fixes the distance to last scattering to be close to the $\Lambda$CDM expectation, which requires to compensate a positive $1+w$ at low redshift with a negative value at higher redshift Linder:2007ka . This adds an additional challenge when trying to interpret the DESI results in terms of physical models. The $w_{0}w_{a}$-model offers a phenomenological description of the equation of state of dark energy at the background level, but it fails to provide a satisfactory description for dark energy perturbations. The dark energy perturbations are usually taken to be those of an canonical minimally-coupled scalar field component, but this model lacks a physical mechanism to safely cross the phantom divide. To avoid singularities in the fluid equation of motion, one commonly employs the parameterized post-Friedman approach to metric perturbations Fang:2008sn . Although this framework is extensively used in the literature, it provides only an effective prescription of the perturbation evolution and is not based on a full physical model. Hence, a physical model for dark energy perturbations is necessary.
@@
Our paper is structured as follows. We describe the analysis procedure and data in Sec. II Our final results are presented in Sec. III. We draw conclusions in Sec. IV.
@@ II.2 Methodology
We perform Markov chain Monte Carlo (MCMC) analyses to sample the posterior distributions using the Metropolis-Hastings algorithm. We adopt a Gelman-Rubin Gelman:1992zz convergence criterion $R-1<0.02$, except in the MG model, where we require a still sufficient threshold of $R-1<0.1$ (unless otherwise stated) due to the higher computational cost. The plots and marginalized constraints are generated with the publicly available getdist package.  [^3] https://getdist.readthedocs.io/en/latest/
In the minimal $\Lambda$CDM model we vary 6 standard cosmological parameters within broad uniform priors: $\omega_{cdm}=\Omega_{c}h^{2}$, $\omega_{b}=\Omega_{b}h^{2}$, $H_{0}$, $\ln(10^{10}A_{s})$, $n_{s}$ and $\tau$. In the $w_{0}w_{a}$ scenario we sample two additional parameters, $(w_{0},w_{a})$, which parameterize a time-varying equation of state for dark energy. We adopt the uniform priors on the dark energy parameters, $w_{0}\in[-3,1]$ and $w_{a}\in[-3,2]$, as specified in DESI:2024mwx . Since we explore the parameter space that allows the dark energy equation of state to cross the $w=-1$ boundary, we use the parameterized post-Friedman approach Fang:2008sn to compute dark energy perturbations in the $w_{0}w_{a}$ scenario. Additionally, we impose the condition $w_{0}+w_{a}<0$, ensuring a matter domination stage in the past. In Horndeski theory, we additionally vary two EFT amplitudes, $c_{B}$ and $c_{M}$, within broad uniform priors while fixing $c_{K}=1$. We also set $c_{T}=0$ to address the tight constraint on the speed of propagation of gravitational waves. In all cases, we assume a spatially flat universe with two massless and one massive neutrino species with $m_{\nu}=0.06\,\text{eV}$. In the $\Lambda$CDM and $w_{0}w_{a}{\rm CDM}$ models, we use the Halofit module to compute the non-linear matter power spectrum, whereas in general Horndeski theory only the linear prediction is utilized.
@@ II.3.1 CMB
Recently, a new Planck data release, PR4, has been made available, involving a reprocessing of the data from both the LFI and HFI channels with the latest and most accurate pipeline, NPIPE Planck:2020olo . This release features lower levels of noise and systematics, as well as slightly more data. In our analysis, we use the official PR3 plik likelihood for the following reasons. First, shifts in parameter estimates inferred from the PR3 and PR4 Planck likelihoods become non-significant after combining with CMB lensing likelihoods from Planck and ACT. Second, we retain plik as our baseline to closely align with the DESI analysis.
@@ II.3.2 BAO
In our baseline analysis we utilize the DESI DR1 BAO data as specified in Tab. I of Ref. DESI:2024mwx , which consists of the BSG sample in the redshift range $0.40.1<z<$ 0.4, LRG1 and LRG2 samples in $0.60.4<z<$ 0.6 and $0.80.6<z<$ 0.8, combined LRG3+ELG1 sample in $1.10.8<z<$ 1.1, ELG2 sample in $1.61.1<z<$ 1.6, quasar sample in $2.10.8<z<$ 2.1 and Lyman-$\alpha$ Forest sample spanning $4.161.77<z<$ 4.16. We will refer to this BAO data as “DESI BAO”.
@@ III.1 $w_{0}w_{a}$ model
Our parameter estimates align closely with the results of the official DESI analysis. The minimum $\Delta\chi^{2}$ value between the $w_{0}w_{a}$ and $\Lambda$CDM models is $-8.8$, indicating a moderate $2.5\sigma$ preference for the $w_{0}w_{a}$ scenario in full agreement with the DESI findings.
@@
Figure 2: $\chi^{2}/N_{\rm data}$ values for the individual BAO measurements computed in the best-fit $w_{0}w_{a}$ model to the $\rm CMB+DESI+SN$ (red) and $\rm CMB+(DESI^{\star}\!+SDSS)+SN$ (blue) data. $N_{\rm data}$ represents the number of data points in a given experiment, as provided in Tab. 1. The BAO data not used in the fitting procedure are displayed with lighter color.
For the full DESI data analysis, the largest contributions to the $\chi^{2}$ come from LRG1 amd LRG2 points, with values of $\chi^{2}=5.6$ and $3.2$, respectively. When using the $\rm(DESI^{\star}\!+SDSS)$ data, we found a reasonable fit to each individual BAO measurements involved in the fitting procedure. Interestingly, this noticeably degrades the fit to the DESI measurements that were not included in the fitting procedure.
@@
Another important observation relates to the LRG2 data point. The ${}^{\star}\!$+SDSS analysis shows a poor fit to the LRG2, with a minimum $\chi^{2}=6.6$. The fit quality improves significantly to $\chi^{2}=3.2$ when the entire DESI data is fitted. This behaviour can be traced to the discrepancy between the LRG2 and eBOSS DR16 results. The level of disagreement between these datasets is close to $3\sigma$, mainly for $D_{M}/r_{d}$ and $D_{V}/r_{d}$ DESI:2024mwx . Given that the eBOSS DR16 is fully consistent with the Planck best-fit $\Lambda$CDM cosmology eBOSS:2020yzd , replacing the DESI BAO measurement at $z_{\rm eff}=0.71$ with the SDSS measurement at $z_{\rm eff}=0.70$ shifts the parameter constraints in the $(w_{0},w_{a})$ plane towards $\Lambda{\rm CDM}$.
@@ III.2 MG model
We now explore a general Horndeski scalar-tensor theory that in principle allows the dark energy equation of state to evolve across the phantom divide, $w=-1$. First of all, we found that our results are not sensitive to the kineticity parameter $\alpha_{K}$. This is expected, since $\alpha_{K}$ has minimal impact on sub-horizon physics Bellini:2014fua ; Noller:2018wyv . For definiteness, we fix $\alpha_{K}=1$ in our analysis.
Fig. 3 shows the posterior distributions in the $w_{0}-w_{a}-\alpha_{B}-\alpha_{M}$ space for the MG model.
Figure 3: 68% and 95% marginalized constraints on $w_{0}$, $w_{a}$, $\alpha_{B}$ and $\alpha_{M}$ from the CMB+DESI+SN data in the MG model.
@@
$&w_{0}=-0.856\pm 0.062,\\ &w_{a}=-0.53_{-0.26}^{+0.28},\\ &\alpha_{B}=0.82_{-0.46}^{+0.31},\\ &\alpha_{M}=0.88_{-1.28}^{+0.75}.\end{aligned}\ \right\}\quad\mbox{$\rm CMB\!+\!DESI\!+\!SN$}$ (3)
@@
The MG model offers a satisfactory physical description of time-evolving dark energy, allowing for the crossing of the phantom divide, $w=-1$. In particular, the $w_{0}$ posterior deviates from $-1$ at the $2.4\sigma$ level. The minimum $\Delta\chi^{2}$ value between the MG model and the concordance cosmology is $-12.2$ for 4 additional degrees of freedom. This reflects a slightly lower preference for evolving dark energy at the $2.4\sigma$ level when compared to the $w_{0}w_{a}{\rm CDM}$ scenario.
Constraints on the MG parameters are primarily driven by the late integrated Sachs-Wolfe (ISW) effect. The ISW effect especially excludes cosmologies with large $\alpha_{B}$ and $\alpha_{M}$ values, as it leads to a large excess of power in the TT CMB power spectrum on large scales Zumalacarregui:2016pph . For not very large $c_{i}$, the effect of modified gravity conspires to prevent large power in the TT CMB power spectrum, resulting in the particular degeneracy directions in the $\alpha_{M}$-$\alpha_{B}$ plane Noller:2018wyv . The data shows a mild preference for positive values of the braiding parameter, $\alpha_{B}>0.11$ at $95\%$ CL., in agreement with the previous study Noller:2018wyv . The lower boundary for $\alpha_{M}$ is driven by the onset of gradient instabilities.
While $\alpha_{M}=0$ provides a good fit to the data, non-minimal coupling play an important role in the MG constraints. For the scalar minimally coupled to gravity ($\alpha_{M}=0$), the region $\alpha_{B}<0.4$ is excluded due to the presence of gradient instabilities. A non-zero $\alpha_{M}$ shifts the gradient stability condition, making cosmologies with small $\alpha_{B}$ viable. The region of small $\alpha_{B}$ is then disfavoured by the data, as it implies too large $\alpha_{M}$, which leads to large power in the $C_{\ell}^{\rm TT}$ via the late-time ISW effect. Thus, non-minimal coupling provides more flexibility, enabling the data to explore the full parameter space in the MG scenario.
It is important to note that measurements of the late-time fluctuation amplitude significantly improves the MG constraints by breaking the degeneracy directions in the $\alpha_{M}$-$\alpha_{B}$ plane, as discussed in Noller:2018wyv . In this work, our primary goal is to demonstrate a proof-of-principle theory that accommodates the crossing of the phantom divide, evidenced by the CMB+DESI+SN data. To streamline the analysis, we focus here on the distance measurements. We leave a detailed analysis of the MG scenario with all available data to future work.
@@ III.3 Comparison And Discussion
The MG scenario shows a slightly weaker preference for evolving dark energy, as evidenced by the $w_{0}$ and $w_{a}$ posteriors shifting towards $\Lambda$CDM. This behaviour is mainly attributed to the CMB degeneracies between the dark energy parameters and $\alpha_{B}$ shown in Fig. 3. However, the results from the MG and $w_{0}w_{a}{\rm CDM}$ models are fully compatible.
@@
We found that non-linear corrections to the matter power spectrum are important when calculating the minimum $\chi^{2}$. Specifically, non-linear effects significantly increase the power of the lensing potential power spectrum at small angular scales, which, if not included, worsens the fit to the CMB lensing data. To ensure a meaningful comparison, we utilize the Halofit prescription for the non-linear matter power spectrum for both models. It is important to note that the Halofit fitting formula was calibrated against $\Lambda{\rm CDM}$ and therefore is not expected to accurately capture the non-linear behavior of the matter power spectrum for modified gravity. Note that the $w_{0}w_{a}{\rm CDM}$ model has two additional parameters compared to $\Lambda{\rm CDM}$, whereas the MG model has four extra parameters.
Both models yield nearly the same goodness-of-fit for the DESI and SN data. The MG scenario shows a better fit to the CMB, but this preference is not statistically significant given the different number of model parameters. Although the MG scenario improves the overall $\chi^{2}$, it shows a similar preference over $\Lambda{\rm CDM}$: $2.4\sigma$ for the MG model versus $2.5\sigma$ for the $w_{0}w_{a}{\rm CDM}$ model. We conclude that the MG model provides a successful physical explanation for DESI’s preference for evolving dark energy.

### v2

@@ Abstract
Taking this preference seriously, we study a general Horndeski scalar-tensor theory, which provides a physical mechanism to safely cross the phantom divide, $w=-1$. Utilizing the Effective Field Theory of dark energy and adopting the $w_{0}w_{a}{\rm CDM}$ background cosmological model, we derive constraints on the parameters $w_{0}=-0.856\pm 0.062$ and $w_{a}=-0.53_{-0.26}^{+0.28}$ at $68\%$ CL. from Planck CMB, Planck and ACT CMB lensing, DESI BAO, and Pantheon+ datasets, showing good consistency with the standard $w_{0}w_{a}{\rm CDM}$ model. The modified gravity model gives results discrepant with $\Lambda$CDM at the $2.4\sigma$ level, while for $w_{0}w_{a}{\rm CDM}$ it is at $2.5\sigma$, based on the best-fit $\chi^{2}$ values. We conclude that modified gravity offers a viable physical explanation for DESI’s preference for evolving dark energy.
@@ I Introduction
One particular aspect of the preference of the DESI+CMB+SN data for an evolving dark energy equation of state is that the preferred models cross the ‘phantom divide’, $w=-1$, a scenario supported by some dark energy models DESI:2024kob and model-agnostic expansion history reconstructions DESI:2024aqx ; Yang:2024kdo . To some extent this is driven by the CMB data, which fixes the distance to last scattering to be close to the $\Lambda$CDM expectation, which requires to compensate a positive $1+w$ at low redshift with a negative value at higher redshift Linder:2007ka . This adds an additional challenge when trying to interpret the DESI results in terms of physical models. The $w_{0}w_{a}$-model offers a phenomenological description of the equation of state of dark energy at the background level, but it fails to provide a satisfactory description for dark energy perturbations. The dark energy perturbations are usually taken to be those of an canonical minimally-coupled scalar field component, but this model lacks a physical mechanism to safely cross the phantom divide. To avoid singularities in the fluid equation of motion, one commonly employs the parameterized post-Friedman approach to metric perturbations Fang:2008sn . Although this framework is extensively used in the literature, it provides only an effective prescription of the perturbation evolution and is not based on a full physical model. Hence, a physical model for dark energy perturbations is necessary.
@@
Our paper is structured as follows. We describe the analysis procedure and data in Sec. II Our final results are presented in Sec. III. We draw conclusions in Sec. IV. Additional parameter constraints are given in Appendix A.
@@ II.2 Methodology
We perform Markov chain Monte Carlo (MCMC) analyses to sample the posterior distributions using the Metropolis-Hastings algorithm. We adopt a Gelman-Rubin Gelman:1992zz convergence criterion $R-1<0.02$ (unless otherwise stated). The plots and marginalized constraints are generated with the publicly available getdist package.  [^3] https://getdist.readthedocs.io/en/latest/
In the minimal $\Lambda$CDM model we vary 6 standard cosmological parameters within broad uniform priors: $\omega_{cdm}=\Omega_{c}h^{2}$, $\omega_{b}=\Omega_{b}h^{2}$, $H_{0}$, $\ln(10^{10}A_{s})$, $n_{s}$ and $\tau$. In the $w_{0}w_{a}$ scenario we sample two additional parameters, $(w_{0},w_{a})$, which parameterize a time-varying equation of state for dark energy. We adopt the uniform priors on the dark energy parameters, $w_{0}\in[-3,1]$ and $w_{a}\in[-3,2]$, as specified in DESI:2024mwx . Since we explore the parameter space that allows the dark energy equation of state to cross the $w=-1$ boundary, we use the parameterized post-Friedman approach Fang:2008sn to compute dark energy perturbations in the $w_{0}w_{a}$ scenario. Additionally, we impose the condition $w_{0}+w_{a}<0$, ensuring a matter domination stage in the past. In Horndeski theory, we additionally vary two EFT amplitudes, $c_{B}$ and $c_{M}$, within broad uniform priors while fixing $c_{K}=1$. We also set $c_{T}=0$ to address the tight constraint on the speed of propagation of gravitational waves. In all cases, we assume a spatially flat universe with two massless and one massive neutrino species with $m_{\nu}=0.06\,\text{eV}$. In the $\Lambda$CDM and $w_{0}w_{a}{\rm CDM}$ models, we use the Halofit module to compute the non-linear matter power spectrum. Since the Halofit fitting formula was not calibrated for modified gravity scenarios, we conservatively adopt the linear prediction within the general Horndeski theory.  [^4] Employing the Halofit module in the MG scenario does not significantly affect the posterior distribution. In particular, the shifts in cosmological parameters w.r.t the baseline analysis do not exceed $0.5\sigma$ in terms of statistical uncertainty, with the modified gravity parameters shifting by less than $0.2\sigma$.
@@ II.3.1 CMB
Recently, a new Planck data release, PR4, has been made available, involving a reprocessing of the data from both the LFI and HFI channels with the latest and most accurate pipeline, NPIPE Planck:2020olo . This release features lower levels of noise and systematics, as well as slightly more data. In our analysis, we use the official PR3 plik likelihood for the following reasons. First, shifts in parameter estimates inferred from the PR3 and PR4 Planck likelihoods become non-significant after combining with CMB lensing likelihoods from Planck and ACT. Second, we retain plik as our baseline to closely align with the DESI analysis DESI:2024mwx .
@@ II.3.2 BAO
In our baseline analysis we utilize the DESI DR1 BAO data as implemented in the official DESI likelihood.  [^6] The DESI likelihood is publicly available at https://github.com/cosmodesi/desilike This dataset includes the BGS sample in the redshift range $0.40.1<z<$ 0.4, LRG1 and LRG2 samples in $0.60.4<z<$ 0.6 and $0.80.6<z<$ 0.8, combined LRG3+ELG1 sample in $1.10.8<z<$ 1.1, ELG2 sample in $1.61.1<z<$ 1.6, quasar sample in $2.10.8<z<$ 2.1 and Lyman-$\alpha$ Forest sample spanning $4.161.77<z<$ 4.16. We will refer to this BAO data as “DESI BAO”.
@@ III.1 $w_{0}w_{a}$ model
Our parameter estimates align closely with the results of the official DESI analysis. The minimum $\Delta\chi^{2}$ value between the $w_{0}w_{a}$ and $\Lambda$CDM models is $-8.8$, indicating a preference for an evolving dark energy equation of state at the $2.4\sigma$ level. 777 The likelihood-ratio test applied here evaluates how likely the observed outcome was under the null hypothesis compared to the alternative. This does not assign probabilities to the models themselves. We interpret the resulting p-value in terms of a $\sigma$-interval to provide a more intuitive result.
@@
Figure 2: $\chi^{2}/N_{\rm data}$ values for the individual BAO measurements computed in the best-fit $w_{0}w_{a}$ model to the $\rm CMB+DESI+SN$ (red) and $\rm CMB+(DESI^{\star}\!+SDSS)+SN$ (blue) data. $N_{\rm data}$ represents the number of data points in a given experiment, as provided in Tab. 1. Results for the BAO data not used in the fitting procedure are displayed with open markers.
For the full DESI data analysis, the largest contributions to the total $\chi^{2}$ come from LRG1 and LRG2 points, with values of $\chi^{2}=5.6$ and $3.2$, respectively. When using the $\rm(DESI^{\star}\!+SDSS)$ data, we found a reasonable fit to each individual BAO measurements involved in the fitting procedure. Interestingly, this noticeably degrades the fit to the DESI measurements that were not included in the fitting procedure.
@@
Another important observation relates to the LRG2 data point. When this measurement is excluded, the $w_{0}w_{a}{\rm CDM}$ model shows a poor fit to the LRG2, with a minimum $\chi^{2}=6.6$. The fit quality improves significantly to $\chi^{2}=3.2$, when the entire DESI dataset is considered. This behaviour can be traced to the discrepancy between the LRG2 and eBOSS DR16 results. The level of disagreement between these datasets is close to $3\sigma$, mainly for $D_{M}/r_{d}$ and $D_{V}/r_{d}$ DESI:2024mwx . Given that the eBOSS DR16 is fully consistent with the Planck best-fit $\Lambda$CDM cosmology eBOSS:2020yzd , replacing the DESI BAO measurement at $z_{\rm eff}=0.71$ with the SDSS measurement at $z_{\rm eff}=0.70$ shifts the parameter constraints in the $(w_{0},w_{a})$ plane towards $\Lambda{\rm CDM}$.
@@ III.2 MG model
We now explore a general Horndeski scalar-tensor theory that in principle allows the dark energy equation of state to evolve across the phantom divide, $w=-1$. First of all, we found that our results are not sensitive to the kineticity parameter $c_{K}$. This is expected, since $c_{K}$ has minimal impact on sub-horizon physics Bellini:2014fua ; Noller:2018wyv . For definiteness, we fix $c_{K}=1$ in our analysis.
Fig. 3 shows the posterior distributions in the $w_{0}-w_{a}-c_{B}-c_{M}$ space for the MG model.
Figure 3: 68% and 95% marginalized constraints on $w_{0}$, $w_{a}$, $c_{B}$ and $c_{M}$ from the CMB+DESI+SN data in the MG model.
@@
$&w_{0}=-0.856\pm 0.062,\\ &w_{a}=-0.53_{-0.26}^{+0.28},\\ &c_{B}=0.82_{-0.46}^{+0.31},\\ &c_{M}=0.88_{-1.28}^{+0.75}.\end{aligned}\ \right\}\quad\mbox{$\rm CMB\!+\!DESI\!+\!SN$}$ (3)
@@
The MG model offers a satisfactory physical description of time-evolving dark energy, allowing for the crossing of the phantom divide, $w=-1$. In particular, the $w_{0}$ posterior deviates from $-1$ at the $2.4\sigma$ level. The minimum $\Delta\chi^{2}$ value between the MG model and the concordance cosmology is $-12.2$ for 4 additional degrees of freedom. This reflects a slightly lower preference for evolving dark energy, when compared to the $w_{0}w_{a}{\rm CDM}$ scenario, at the $2.4\sigma$ level relative to $\Lambda{\rm CDM}$.
Constraints on the MG parameters are primarily driven by the late integrated Sachs-Wolfe (ISW) effect. The ISW effect especially excludes cosmologies with large $c_{B}$ and $c_{M}$ values, as it leads to a large excess of power in the TT CMB power spectrum on large scales Zumalacarregui:2016pph . For not very large $c_{i}$, the effect of modified gravity conspires to prevent large power in the TT CMB power spectrum, resulting in the particular degeneracy directions in the $c_{M}$-$c_{B}$ plane Noller:2018wyv . The data shows a mild preference for positive values of the braiding parameter, $c_{B}>0.11$ at $95\%$ CL., in agreement with the previous study Noller:2018wyv . The lower boundary for $c_{M}$ is driven by the onset of gradient instabilities.
While $c_{M}=0$ provides a good fit to the data, non-minimal coupling play an important role in the MG constraints.  [^9] Ref. Wolf:2024stt highlights the importance of non-minimal gravity coupling in explaining DESI’s preference for evolving dark energy. For the scalar minimally coupled to gravity ($c_{M}=0$), the region $c_{B}<0.4$ is excluded due to the presence of gradient instabilities. A non-zero $c_{M}$ shifts the gradient stability condition, making cosmologies with small $c_{B}$ viable. The region of small $c_{B}$ is then disfavoured by the data, as it implies too large $c_{M}$, which leads to large power in the $C_{\ell}^{\rm TT}$ via the late-time ISW effect. Thus, non-minimal coupling provides more flexibility, enabling the data to explore the full parameter space in the MG scenario.
It is important to note that measurements of the late-time fluctuation amplitude significantly improves the MG constraints by breaking the degeneracy directions in the $c_{M}$-$c_{B}$ plane, as discussed in Noller:2018wyv . In this work, our primary goal is to demonstrate a proof-of-principle theory that accommodates the crossing of the phantom divide, evidenced by the CMB+DESI+SN data. To streamline the analysis, we focus here on the distance measurements. We leave a detailed analysis of the MG scenario with all available data to future work.
@@ III.3 Comparison And Discussion
The MG scenario shows a slightly weaker preference for evolving dark energy, as evidenced by the $w_{0}$ and $w_{a}$ posteriors shifting towards $\Lambda$CDM. This behaviour is mainly attributed to the CMB degeneracies between the dark energy parameters and $c_{B}$ shown in Fig. 3. However, the results from the MG and $w_{0}w_{a}{\rm CDM}$ models are fully compatible.
@@
We found that non-linear corrections to the matter power spectrum are important when calculating the minimum $\chi^{2}$. Specifically, non-linear effects significantly increase the power of the lensing potential power spectrum at small angular scales, which, if not included, worsens the fit to the CMB lensing data. To ensure a meaningful comparison, we utilize the Halofit prescription Bird:2011rb for the non-linear matter power spectrum for both models. It is important to note that the Halofit fitting formula was calibrated against $\Lambda{\rm CDM}$ and therefore is not expected to accurately capture the non-linear behavior of the matter power spectrum for modified gravity. Note that the $w_{0}w_{a}{\rm CDM}$ model has two additional parameters compared to $\Lambda{\rm CDM}$, whereas the MG model has four extra parameters.
Both models yield nearly the same goodness-of-fit for the DESI and SN data. The MG scenario shows a better fit to the CMB, but this preference is not statistically significant given the different number of model parameters. Overall, the data shows a similar level of preference for both scenarios relative to $\Lambda{\rm CDM}$ based on the best-fit $\chi^{2}$ values: $2.4\sigma$ for the MG model and $2.5\sigma$ for the $w_{0}w_{a}{\rm CDM}$ model.  [^10] Ref. Wang:2024hks reported a similar level of preference for modified gravity from the CMB+DESI data when using the $(\mu,\eta)$ parameterization for linear perturbations.
Additionally, we evaluate the preference for the non-standard scenarios relative to $\Lambda$CDM in the context of a model-selection analysis. First, we apply the Akaike Information Criteria (AIC) 1100705 , defined by ${\rm AIC}=\chi^{2}_{\rm min}+2N_{p}$, where $N_{p}$ represents the number of free parameters in the model. We found similar preferences for the $w_{0}w_{a}{\rm CDM}$ and MG models relative to $\Lambda$CDM: $\Delta{\rm AIC}=-4.8$ and $-4.2$, respectively. Next, we employ the Deviance Information Criterion (DIC) Liddle:2007fy , defined as ${\rm DIC}=\chi^{2}(\theta_{p})+2p_{D}$, where $\theta_{p}$ represents the parameters at the maximum point of the posterior, and $p_{D}=\langle\chi^{2}\rangle-\chi^{2}(\theta_{p})$ is the effective number of parameters that the data constrain, where the average is over the posterior Raveri:2018wln . We obtain $\Delta{\rm DIC}=-4.9$ and $+1.7$ for the $w_{0}w_{a}{\rm CDM}$ and MG models over $\Lambda{\rm CDM}$. These results suggest a moderate preference for the $w_{0}w_{a}{\rm CDM}$ scenario, while the MG model shows no significant difference from $\Lambda{\rm CDM}$, according to the Jeffreys’ scale Spiegelhalter:2002yvw .
In conclusion, while the MG scenario significantly improves the fit quality, it does not show a preference over $\Lambda{\rm CDM}$. The MG model still offers a physical mechanism to safely cross the phantom divide and is capable of accommodate DESI’s preference for evolving dark energy.